# Bose-Einstein condensation in exotic lattice geometries

### v1

@@ IV.2 Field-theoretic description of the Mott transition
For further analysis, we assume that $J_{ij}$ takes values zero or $J>0$, and the eigenvalues of $J_{ij}/J$ shall be denoted by $\lambda_{k}$. In this notation, the tight-binding energy spectrum is given by $\epsilon_{k}=-J\lambda_{k}$. With this, $M_{kk}=\frac{1}{g(0)}-J\lambda_{k}=\frac{1}{g(0)}+\epsilon_{k}$. Let us now focus on the parameter regime $10<\mu/U<$ 1, where $n_{0}=1$, and $g(0)=\frac{\mu+U}{\mu(U-\mu)}>0$, and first consider a regular lattice where $J_{ij}/J$ can be diagonalized via Fourier transform, yielding an energy band with a lower band edge at $\epsilon_{\rm min}=-2dJ<0$, in the case of a $d$-dimensional hypercubic lattice. Thus, the matrix $M$ becomes singular when $\frac{1}{g(0)}-2dJ=0$. This means the system enters the superfluid regime for $J>\frac{\mu(U-\mu)}{4(\mu+U)}$. This expression reproduces the mean-field value of the first Mott lobe.
@@
where $-J\lambda_{k}$ denotes an eigenstate of the tight-binding matrix. Upon increasing the hopping strength $J>0$, this condition is first met for the ground state (i.e. $k=0$) at a critical hopping strength $J_{\rm crit}=-\frac{1}{g(0)\lambda_{0}}$, as also the case for the regular lattice. The ground state energy of the Sierpiński gasket is identical to the one of a regular square lattice, $\epsilon_{0}=-4J$, and hence the two systems have identical Mott lobes. In the case of regular lattices, $\epsilon_{0}=-zJ$ is directly given by the coordination number $z$ of the lattice, and it immediately follows that in the case of a triangular lattice ($z=6$), the height of the lobe is suppressed. Specifically, the quantitative comparison of a Sierpiński gasket and a regular triangle shows an increased Mott phase for the gasket, and the heights at the tips of the lobes takes a ratio $3/2$.
@@ V Conclusions
Geometry and dimensionality can strongly influence the behavior of a quantum system. Many studies [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77] have been devoted to the effect of fractal geometry on the single-particle level. In the present paper, we have explored the role of exotic geometries on quantum-statistical and quantum many-body behavior. Specifically, for fractal and hyperbolic lattices, we have scrutinized the phenomenon of Bose-Einstein condensation of an ideal gas, as well as the zero-temperature Mott-insulator–to-superfluid transition in the presence of strong on-site repulsion. For all fractal geometries considered, we found that the condensation temperature is dramatically lowered as compared to regular lattices of similar size, but the thermal depletion as a function of temperature has a different functional behavior, with a heavy tail above the critical temperature. Condensate fluctuations are found to be enlarged by the fractal geometries, with an anomalous scaling exponent $>2$. As in regular lattices in less than three dimensions, the critical temperature in fractal lattices decreases with system size and is expected to be zero in thermodynamically large systems. A strikingly different behavior has been seen in hyperbolic lattices: Despite the fact that these lattices can be embedded in two spatial dimensions, the critical temperature increases with system size. In the presence of strong interactions, we observed changes in the shape of the Mott lobe, obtained from a cluster Gutzwiller ansatz due to fractal geometry. In accordance with the fractal dimension of the lattice, the shape appears to be intermediate to the well-known cases of 1D and 2D Mott lobes. Our theoretical research is aligned with current experimental trends which have started to realize the topologies analyzed here in state-of-the-art quantum simulators, including photonic wave-guide arrays and tweezer-assembled Rydberg-atom arrays. We believe that our results will spur systematic experimental exploration of quantum matter in non-Euclidean settings, where geometry itself becomes a tunable control parameter that can unveil genuinely new phases and critical behavior.
Appendix A Effective action description of the Mott phase
@@
[…]
Here, we shall briefly comment on some details in the real-time/Keldysh formalism. As mentioned earlier, the temporal integral shall enclose a contour with forward and backward paths along the real-time axis. However, it is convenient to write the time-contour integral as a single forward integral (from $\infty$- ∞ to $\infty$), and equip each temporal degree of freedom with an additonal path index in order to differentiate between forward and backward path. With this, the fields are doubled into two-component vector fields, and the Green function $G_{ij}(t_{1},t_{2})$ becomes a $2\times 2$ matrix. At this point, it is very convenient to make a rotation within this path-index space, known as Keldysh rotation. This rotaion makes one element in the Green function matrix vanish, and the off-diagonal elements of this matrix are the retarded and advanced Green functions. Within this Keldysh space, the two-component fields have a "classical" field component, $\Psi_{i,{\rm cl}}(t),\Psi_{i,{\rm cl}}(t)^{*}$, which is the sum of the fields on the forward and backward path, and a "quantum" field component, $\Psi_{i,{\rm q}}(t),\Psi_{i,{\rm q}}(t)^{*}$, which is the difference of the fields on the forward and backward path.
@@
[…]
@@
In order to find the inverse Green function, $[G_{ij}^{\rm(ret)}(\omega)]^{-1}$, we use a hopping expansion in which the unperturbed Hamiltonian are all the local terms of the Bose-Hubbard model. The Green functions for the local Hamiltonian are easily obtained. For instance, the retarded Green function in this local limit, $G_{ij}^{(0,{\rm ret})}(\omega)\equiv G_{i}^{(0,{\rm ret})}(\omega)\delta_{ij}$ reads
@@ Acknowledgements.
T.G. acknowledges the financial support received from the IKUR Strategy under the collaboration agreement between the Ikerbasque Foundation and DIPC on behalf of the Department of Education of the Basque Government, as well as funding by the Department of Education of the Basque Government through the project PIBA_2023_1_0021 (TENINT), by the Agencia Estatal de Investigación (AEI) through Proyectos de Generación de Conocimiento PID2022-142308NA-I00 (EXQUSMI), and that this work has been produced with the support of a 2023 Leonardo Grant for Researchers in Physics, BBVA Foundation. The BBVA Foundation is not responsible for the opinions, comments, and contents included in the project and/or the results derived therefrom, which are the total and absolute responsibility of the authors. ICFO-QOT group acknowledges support from: European Research Council AdG NOQIA; MCIN/AEI (PGC2018-0910.13039/501100011033, CEX2019-000910-S/10.13039/501100011033, Plan National FIDEUA PID2019-106901GB-I00, Plan National STAMEENA PID2022-139099NB, I00, project funded by MCIN/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR" (PRTR-C17.I1), FPI); QUANTERA DYNAMITE PCI2022-132919, QuantERA II Programme co-funded by European Union’s Horizon 2020 program under Grant Agreement No 101017733; Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda; Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program; Barcelona Supercomputing Center MareNostrum (FI-2023-3-0024); Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, European Commission, European Climate, Infrastructure and Environment Executive Agency (CINEA), or any other granting authority. Neither the European Union nor any granting authority can be held responsible for them (HORIZON-CL4-2022-QUANTUM-02-SGA PASQuanS2.1, 101113690, EU Horizon 2020 FET-OPEN OPTOlogic, Grant No 899794, QU-ATTO, 101168628), EU Horizon Europe Program (This project has received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101080086 NeQSTGrant Agreement 101080086 — NeQST); ICFO Internal “QuantumGaudi” project. B.J.-D. and A.R.-F. funding from Grant PID2023-147475NB-I00 funded by MICIU/AEI/10.13039/501100011033 and FEDER, UE, by grants 2021SGR01095 from Generalitat de Catalunya, and by Project CEX2019-000918-M of ICCUB (Unidad de Excelencia María de Maeztu).

### v2

@@ IV.2 Field-theoretic description of the Mott transition
For further analysis, we assume that $J_{ij}$ takes values zero or $J>0$, and the eigenvalues of $J_{ij}/J$ shall be denoted by $\lambda_{k}$. In this notation, the tight-binding energy spectrum is given by $\epsilon_{k}=-J\lambda_{k}$. With this, $M_{kk}=\frac{1}{g(0)}-J\lambda_{k}=\frac{1}{g(0)}+\epsilon_{k}$. Let us now focus on the parameter regime $0<\mu/U<1$, where $n_{0}=1$, and $g(0)=\frac{\mu+U}{\mu(U-\mu)}>0$, and first consider a regular lattice where $J_{ij}/J$ can be diagonalized via Fourier transform, yielding an energy band with a lower band edge at $\epsilon_{\rm min}=-2dJ<0$, in the case of a $d$-dimensional hypercubic lattice. Thus, the matrix $M$ becomes singular when $\frac{1}{g(0)}-2dJ=0$. This means the system enters the superfluid regime for $J>\frac{\mu(U-\mu)}{4(\mu+U)}$. This expression reproduces the mean-field value of the first Mott lobe.
@@
where $J\lambda_{k}$ denotes an eigenstate of the tight-binding matrix. Upon increasing the hopping strength $J>0$, this condition is first met for the ground state (i.e. $k=0$) at a critical hopping strength $J_{\rm crit}=-\frac{1}{g(0)\lambda_{0}}$, as also the case for the regular lattice. The ground state energy of the Sierpiński gasket is identical to the one of a regular square lattice, $\epsilon_{0}=-4J$, and hence the two systems have identical Mott lobes. In the case of regular lattices, $\epsilon_{0}=-zJ$ is directly given by the coordination number $z$ of the lattice, and it immediately follows that in the case of a triangular lattice ($z=6$), the height of the lobe is suppressed. Specifically, the quantitative comparison of a Sierpiński gasket and a regular triangle shows an increased Mott phase for the gasket, and the heights at the tips of the lobes takes a ratio $3/2$.
@@ V Conclusions
Geometry and dimensionality can strongly influence the behavior of a quantum system. Many studies [59, 60, 61, 62, 63, 44, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77] have been devoted to the effect of fractal geometry on the single-particle level. In the present paper, we have explored the role of exotic geometries on quantum-statistical and quantum many-body behavior. Specifically, for fractal and hyperbolic lattices, we have scrutinized the phenomenon of Bose‑Einstein condensation of an ideal gas, as well as the zero‑temperature Mott‑insulator-to‑superfluid transition in the presence of strong on-site repulsion. For all fractal geometries considered, we found that the condensation temperature is dramatically lowered as compared to regular lattices of similar size, but the thermal depletion as a function of temperature has a different functional behavior, with a heavy tail above the critical temperature. Condensate fluctuations are found to be enlarged by the fractal geometries, with an anomalous scaling exponent >2>2. As in regular lattices in less than three dimensions, the critical temperature in fractal lattices decreases with system size and is expected to be zero in thermodynamically large systems. A strikingly different behavior has been seen in hyperbolic lattices: Despite the fact that these lattices can be embedded in two spatial dimensions, the critical temperature increases with system size. In the presence of strong interactions, we observed changes in the shape of the Mott lobe, obtained from a cluster Gutzwiller ansatz due to fractal geometry. In accordance with the fractal dimension of the lattice, the shape appears to be intermediate to the well-known cases of 1D and 2D Mott lobes. Our theoretical research is aligned with current experimental trends which have started to realize the topologies analyzed here in state‑of‑the‑art quantum simulators, including photonic wave‑guide arrays and tweezer‑assembled Rydberg‑atom arrays. We believe that our results will spur systematic experimental exploration of quantum matter in non‑Euclidean settings, where geometry itself becomes a tunable control parameter that can unveil genuinely new phases and critical behavior.
Appendix A Cluster Gutzwiller method with basis truncation
Figure 13: Panels (a,b,e,f) show obtained phase diagrams for various clusters, where the black regions correspond to the order parameter $\phi=0$, yellow regions correspond to regions where $|\phi|>0$ and orange regions correspond to regions where $\phi=0$, but $|n/V_{\cal{C}}-1|>1/V_{\cal{C}}$. Black regions correspond to the Mott insulator (MI) phase, while yellow and orange regions indicate the superfluid (SF) phase. Panels (c,d) show the expected number of bosons per site $n/V_{\cal{C}}\coloneqq\sum_{i}\langle GS|\hat{n}_{i}/V_{\cal C}|GS\rangle$. A 1D cluster with $V_{\cal{C}}=9$ sites is used for panels (a-c), a 1D cluster with $V_{\cal{C}}=15$ sites (same as in Figure 11(a)) is used for panels (d,e) and for panel (f), a triangular cluster connected to the bulk only through two couplings at each corner (as shown in the top right). In panel (a), the basis is constructed with $\Delta N_{\max}=1$ and $f_{\max}=5$, while in panels (b,c) $\Delta N_{\max}=V_{\cal{C}}=9$ and $f_{\max}=V_{\cal{C}}=9$. Panels (d-f) use $\Delta N_{\max}=5$ and $f_{\max}=5$, same as in Figures 10 and 11.
In this section, we explain the details of the cluster Gutzwiller method adapted to the fractal geometry.
[…]
We consider fluctuations of the number of bosons around the unit filling, $N_{i}=V_{\cal C}$. Consequently, the composite Hilbert space contains $N_{i}=V_{\cal C}-\Delta N_{\max},\dots,V_{\cal C},\dots,V_{\cal C}+\Delta N_{\max}$, with $\Delta N_{\max}=5$. Next, we truncate the basis, following the scheme proposed by Lühmann [130], i.e. we discard Fock states with (i) large local fluctuations, where the number of particles $n_{i}$ on any site $i$ differs significantly from $1$: $\exists_{i}~|n_{i}-1|>3$ or (ii) large global fluctuations, where the sum over the sites $\sum_{i}|n_{i}-1|>f_{\max}$, with $f_{\max}=5$. We note that this truncation is most accurate inside and near the MI phase, where the ground states have low fluctuations the of occupation operator. When the parameter $\Delta N_{\max}$ is too small, the order parameter $\phi$ may vanish even in the SF phase, when the ground state populates states with $N=V\pm\Delta N_{\max}$ particles. We use $\Delta N_{\max}=5$, which makes the size of the basis limited primarily by the maximum number of global fluctuations $f_{\max}=5$. This is sufficient for determining the Mott lobe boundaries in clusters with $V_{\cal C}=15$ sites, although in the lower-dimensional geometries like the 1D chain and the Sierpiński triangle, we still observe regions in the SF phase where $\phi=0$ (see Fig. 13 (e,f)). We can identify such misclassified regions by checking where the number of particles per site $n/V_{\cal{C}}\coloneqq\sum_{i}\langle GS|\hat{n}_{i}/V_{\cal C}|GS\rangle$ in the ground state differs significantly from $1$: $|n/V_{\cal{C}}-1|>1/V_{\cal{C}}$.
This is illustrated through Figures 13 (a-c), which show results for a 1D cluster with $V_{\cal C}=9$ sites. When $\Delta N_{\max}=1$ and $f_{\max}=5$ (Fig. 13 (a)), orange regions appear that are absent when $\Delta N_{\max}=9$ and $f_{\max}=9$ (Fig. 13 (b)), but the shape of the Mott lobe remains the same. Moreover, we can still distinguish between the MI and SF phases in regions where $\phi=0$ by examining the average number of particles per site $n/V_{\cal{C}}$ (Fig. 13 (c)). The same approach is illustrated for a larger 1D cluster with $V_{\cal C}=15$ sites in Figure 13 (d,e) and also used in generating Figures 10 and 11.
Lastly, it is worth noting that both the internal structure of a cluster and its couplings to the mean-field contribute to the shape of the Mott lobe. The phase diagram in Figure 13 (f) shows how a cluster with the same triangular internal structure as in Figure 11(c), but with couplings matching those of the Sierpiński triangle in Figure 11(b), more closely resembles the latter. At the same time, the Mott lobes of 1D clusters with $V_{\cal{C}}=9$ and $V_{\cal{C}}=15$ sites differ significantly (see Fig. 13 (b,e)), despite being coupled to the mean-field in the same way.
Appendix B Effective action description of the Mott phase
@@
$\Gamma[\Psi,\Psi^{*}]-\hbar^{2}\sum_{i,j}\int dt_{1}\int dt_{2}\Psi_{i}(t_{1})^{*}G_{ij}^{-1}(t_{1},t_{2})\Psi_{j}(t_{2})$ (17)
Here, we shall briefly comment on some details in the real-time/Keldysh formalism. As mentioned earlier, the temporal integral shall enclose a contour with forward and backward paths along the real-time axis. However, it is convenient to write the time-contour integral as a single forward integral (from -$\infty$ to $\infty$), and equip each temporal degree of freedom with an additonal path index in order to differentiate between forward and backward path. With this, the fields are doubled into two-component vector fields, and the Green function $G_{ij}(t_{1},t_{2})$ becomes a $2\times 2$ matrix. At this point, it is very convenient to make a rotation within this path-index space, known as Keldysh rotation. This rotaion makes one element in the Green function matrix vanish, and the off-diagonal elements of this matrix are the retarded and advanced Green functions. Within this Keldysh space, the two-component fields have a "classical" field component, $\Psi_{i,{\rm cl}}(t),\Psi_{i,{\rm cl}}(t)^{*}$, which is the sum of the fields on the forward and backward path, and a "quantum" field component, $\Psi_{i,{\rm q}}(t),\Psi_{i,{\rm q}}(t)^{*}$, which is the difference of the fields on the forward and backward path.
@@
$\Gamma[\Psi,\Psi^{*}]-\hbar^{2}\sum_{i,j}\int d\omega\int dt_{2}\Psi_{i}(\omega)^{*}[G_{ij}(\omega)]^{-1}\Psi_{j}(\omega).$ (18)
@@
In order to find the inverse Green function, $G_{ij}^{\rm(ret)}(\omega)]^{-1}$, we use a hopping expansion in which the unperturbed Hamiltonian are all the local terms of the Bose-Hubbard model. The Green functions for the local Hamiltonian are easily obtained. For instance, the retarded Green function in this local limit, $G_{ij}^{(0,{\rm ret})}(\omega)\equiv G_{i}^{(0,{\rm ret})}(\omega)\delta_{ij}$ reads
@@ Acknowledgements.
T.G. acknowledges the financial support received from the IKUR Strategy under the collaboration agreement between the Ikerbasque Foundation and DIPC on behalf of the Department of Education of the Basque Government, as well as funding by the Department of Education of the Basque Government through the project PIBA_2023_1_0021 (TENINT), by the Agencia Estatal de Investigación (AEI) through Proyectos de Generación de Conocimiento PID2022-142308NA-I00 (EXQUSMI), and that this work has been produced with the support of a 2023 Leonardo Grant for Researchers in Physics, BBVA Foundation. The BBVA Foundation is not responsible for the opinions, comments, and contents included in the project and/or the results derived therefrom, which are the total and absolute responsibility of the authors. ICFO-QOT group acknowledges support from: European Research Council AdG NOQIA; MCIN/AEI (PGC2018-0910.13039/501100011033, CEX2019-000910-S/10.13039/501100011033, Plan National FIDEUA PID2019-106901GB-I00, Plan National STAMEENA PID2022-139099NB, I00, project funded by MCIN/AEI/10.13039/501100011033 and by the “European Union NextGenerationEU/PRTR" (PRTR-C17.I1), FPI); QUANTERA DYNAMITE PCI2022-132919, QuantERA II Programme co-funded by European Union’s Horizon 2020 program under Grant Agreement No 101017733; Ministry for Digital Transformation and of Civil Service of the Spanish Government through the QUANTUM ENIA project call - Quantum Spain project, and by the European Union through the Recovery, Transformation and Resilience Plan - NextGenerationEU within the framework of the Digital Spain 2026 Agenda; Fundació Cellex; Fundació Mir-Puig; Generalitat de Catalunya (European Social Fund FEDER and CERCA program; Barcelona Supercomputing Center MareNostrum (FI-2023-3-0024); Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, European Commission, European Climate, Infrastructure and Environment Executive Agency (CINEA), or any other granting authority. Neither the European Union nor any granting authority can be held responsible for them (HORIZON-CL4-2022-QUANTUM-02-SGA PASQuanS2.1, 101113690, EU Horizon 2020 FET-OPEN OPTOlogic, Grant No 899794, QU-ATTO, 101168628), EU Horizon Europe Program (This project has received funding from the European Union’s Horizon Europe research and innovation program under grant agreement No 101080086 NeQSTGrant Agreement 101080086 — NeQST); ICFO Internal “QuantumGaudi” project. B.J.-D. and A.R.-F. funding from Grant PID2023-147475NB-I00 funded by MICIU/AEI/10.13039/501100011033 and FEDER, UE, by grants 2021SGR01095 from Generalitat de Catalunya, and by Project CEX2019-000918-M of ICCUB (Unidad de Excelencia María de Maeztu). K.D. acknowledges the support of the "la Caixa" Foundation. The project that gave rise to these results received the support of a fellowship from the ”la Caixa” Foundation (ID 100010434). The fellowship code is LCF/BQ/DFI25/13000067. The data that support the findings of this article are openly available [144].